Calculate the ratio of imidazole to the imidazolium ion in a solution at pH 7.4.
Answer #1
Concepts and reason

Acid dissociation constant:

Acid dissociation constant of an acid indicates the extent to which the acid dissociates into ion and a conjugate base. Strong acids will have greater dissociation constants, and hence the value will be greater.

Power of hydrogen ions (pH):

The pH of a solution is defined as the negative logarithmic function of the concentration of hydrogen () ions in an aqueous solution at a particular temperature.

Fundamentals

The Henderson-Hasselbalch equation is shown as follows.

[A ]
pH = PK + log
[HA]

Here,

The power of in a solution is .

The power of acid dissociation constant of a weak acid is .

The concentration of a weak monoprotic acid is .

The concentration of its conjugate base is .

The given of the solution is .

The acid dissociation reaction of imidazolium ion is given below.

imidazolium ion
water
hydronium ion
imidazole
pka = 7.0
н
HN
HN
NH
HOH
н:9-н
acid
base
conjugate acid
conjugate base

Use the Henderson-Hasselbalch equation to calculate the ratio as follows.

[imidazole]
pH =pK, +log;
Simidazolium ion]

Substitute for and for

[imidazole)
7.4 = 7.0+log
[imidazolium ion]
[imidazole)
= 7.4 - 7.0
[imidazolium ion]
[imidazole)
=1004
[imidazolium ion]
[im

So, the ratio of imidazole to the imidazolium ion is .

Ans:

The ratio of imidazole to the imidazolium ion in a solution at of is .